### v1

# KompeteAI: Accelerated Autonomous Multi-Agent System for End-to-End Pipeline Generation for Machine Learning Problems

Abstract
Recent Large Language Model (LLM)-based AutoML systems demonstrate impressive capabilities but face significant limitations such as constrained exploration strategies and a severe execution bottleneck. Exploration is hindered by one-shot methods lacking diversity and Monte Carlo Tree Search (MCTS) approaches that fail to recombine strong partial solutions. The execution bottleneck arises from lengthy code validation cycles that stifle iterative refinement. To overcome these challenges, we introduce KompeteAI, a novel AutoML framework with dynamic solution space exploration. Unlike previous MCTS methods that treat ideas in isolation, KompeteAI introduces a merging stage that composes top candidates. We further expand the hypothesis space by integrating Retrieval-Augmented Generation (RAG), sourcing ideas from Kaggle notebooks and arXiv papers to incorporate real-world strategies. KompeteAI also addresses the execution bottleneck via a predictive scoring model and an accelerated debugging method, assessing solution potential using early stage metrics to avoid costly full-code execution. This approach accelerates pipeline evaluation 6.9 times. KompeteAI outperforms leading methods (e.g., RD-agent, AIDE, and Ml-Master) by an average of 3% on the primary AutoML benchmark, MLE-Bench. Additionally, we propose Kompete-bench to address limitations in MLE-Bench, where KompeteAI also achieves state-of-the-art results.
1 Introduction
Recent research has shifted towards the use of Large Language Models (LLM) as the core reasoning engine for AutoML frameworks, enabling the autonomous generation and testing of end-to-end pipelines that adapt to specific tasks (Li et al. 2024; Jiang et al. 2025; Chi et al. 2024; Liu et al. 2025; Trirat, Jeong, and Hwang 2024; Grosnit et al. 2024; Yang et al. 2025). However, these approaches have critical limitations. Initial ”one-shot” generation can yield diverse ideas but lacks iterative refinement, so a single flawed component - like suboptimal feature engineering — can undermine the entire pipeline without a way to correct it.
More adaptive frameworks based on Monte Carlo Tree Search (MCTS) address this by exploring a tree of potential solutions, but often use constrained exploration and struggle to recombine promising ideas from different branches. Even recent frameworks that combine exploratory search with LLM-based reasoning (Liu et al. 2025), despite achieving state-of-the-art results on MLE-Bench (Chan et al. 2024), are limited by architectural constraints. They may not systematically preserve or merge valuable insights from distant high-performing branches. As a result, valuable solution components can be prematurely discarded.
Even an ideal internal recombination can only reshuffles known ideas. Retrieval‑Augmented Generation (RAG) breaks this limit by injecting external, domain-specific knowledge — enabling the agent to explore solutions beyond its pretrained hypothesis space. This potential remains largely untapped in current implementations. Most systems apply RAG only during early stages; as the problem shifts — from statistical feature engineering to physics-based simulation or chemistry-driven descriptors — the agent cannot retrieve fresh, relevant knowledge, leading to knowledge decay as the pipeline evolves.
Beyond search and knowledge limitations, all frameworks suffer from a severe execution bottleneck. Validating a single solution requires full code execution, which can take hours for complex models. This issue becomes even more problematic during debugging, as errors late in the pipeline force a complete retraining for every fix. As a result, the slow feedback loop discourages major changes and presents a serious scalability challenge.
In this work, we introduce KompeteAI, an autonomous multi-agent framework for structured, multistage pipeline generation. The key innovations compared to prior approaches is presented in Table 1. To efficiently explore and exploit the search space, we employ two core operators. Adding, which dynamically generates novel stage-specific ideas by querying external knowledge sources via an adaptive RAG module, and merging, which intelligently combines the most successful solutions. We address execution bottlenecks through a predictive scoring model that prunes weak solutions early, alongside an accelerated debugging paradigm using simplified code and smaller data samples, dramatically shortening the feedback loop.
Our experiments show that KompeteAI sets a new state-of-the-art on MLE-Bench, outperforming prior methods by an average of 3%, while accelerating the average test-time performance by a factor of 6.9.
Finally, the only publicly available competitive benchmark MLE-Bench suffers from two key limitations. It is excessively large, constructs its test sets by partitioning the original training data, and then compares the resulting scores to the private leaderboard positions that cause evaluation bias. To address these issues, we introduce a new benchmark — Kompete-bench.
Our primary contributions:
  • Stage-Decomposed Multi-Agent Architecture: A state-of-the-art framework that partitions the ML workflow into discrete stages, enabling agents to specialize in focused tasks, dynamically integrate external knowledge sources to enhance exploration diversity, and systematically recombine optimal partial solutions through novel addition and merging operators.
  • An Accelerated Evaluation and Debugging Paradigm: A two-part solution to the execution bottleneck, combining a predictive scoring model and a rapid debugging framework to drastically reduce validation time.
  • A New Benchmark: Kompete-bench — A curated benchmark of recent real-world problems designed to more rigorously evaluate a model’s genuine problem-solving ability, minimizing the influence of memorization or prior exposure.
2 Related Work
Classic AutoML
Classic AutoML frameworks - TPOT, AutoGluon, AutoKeras, LightAutoML, and others (Olson and Moore 2016; Erickson et al. 2020; Jin, Song, and Hu 2019; Vakhrushev et al. 2021; Feurer et al. 2022; LeDell and Poirier 2020; Thornton et al. 2013) — automate data preprocessing, model selection, and hyperparameter tuning via heuristic search, ensembling, and Bayesian optimization. Despite their effectiveness, these systems operate within static search spaces, require manual data preparation and adaptation for each new task, and lack the dynamic coordination and continual learning capabilities inherent to multi‑agent AutoML architectures.
AutoML based LLM
LLM-based AutoML systems have progressed rapidly, offering increasingly autonomous capabilities through dynamic coordination and iterative planning (Li et al. 2024; Chi et al. 2024; Jiang et al. 2025; Trirat, Jeong, and Hwang 2024; Yang et al. 2025; Liu et al. 2025). To analyze and compare these systems meaningfully, we focus on four key aspects that consistently shape their performance and flexibility: exploration strategy, RAG, merge methods, and debugging techniques. These components address critical challenges related to how systems explore the design space, incorporate external knowledge, manage idea diversity, and implement solutions both efficiently and reliably. Notably, RD-Agent’s merge approach amounts to simple, uncontrolled recombination driven by LLM rather than a structured merge algorithm, which results in incoherent or suboptimal pipeline integrations. A summary of how existing systems approach these dimensions is provided in Table 1.
Table 1: Comparison of LLM‑based AutoML systems. Phase‑wise chain: sequential, stage‑by‑stage pipeline construction; without context: searches that ignore information from parallel or prior branches; Multi‑stage expansion: KompeteAI’s staged node growth at every pipeline level; R&D‑phase: retrieval only during initial planning, versus Dynamic RAG: on‑the‑fly literature/web queries; Incremental debugging: iterative draft–debug–refine cycles, versus Multistage debugging: KompeteAI’s multi‑agent, stage‑based error checking; Simple recombination: implicit LLM‑driven fusion of ideas, versus Controlled merger: KompeteAI’s explicit algorithmic branch combination.
Method Exploration Strategy RAG system Debugging method Merge Method
AutoKaggle Phase-wise chain Three-stage phase & Unit tests
SELA MCTS (UCT) without context
AIDE MCTS (UCT) without context Incremental
AutoML-Agent Retrieval-augmented planning R&D-phase Incremental
R&D-Agent Multi-trace exploration R&D-phase Incremental Simple recombination
ML-Master Controlled MCTS with memory Incremental
KompeteAI Multi-stage expansion Dynamic Multistage Controlled merger
Scoring Model
Our approach to model scoring is inspired by similar performance prediction methods developed in Neural Architecture Search (Elsken, Metzen, and Hutter 2019). In NAS, a common approach is to use weight-sharing supernets, where multiple architectures are jointly trained by sharing parameters with a large model, and performance is estimated by evaluating sampled architectures on a validation set (Jawahar et al. 2023a). This method, however, faces challenges such as weight co-adaptation (Bender et al. 2018), capacity bottlenecks (Jawahar et al. 2023b), and gradient conflicts (Gong and Wang 2022). More recently (Jawahar et al. 2023a), demonstrated that LLMs can serve as effective performance predictors in NAS tasks, providing a promising alternative to traditional methods. In contrast to NAS, the AutoML setting typically involves a much broader and less constrained search space, which motivates the exploration of LLM-based performance prediction beyond neural architectures.
Benchmarks
Using Kaggle competitions to evaluate autonomous ML systems has become a modern approach in a number of recent benchmarks due to clear metrics, variety of tasks and the ability to compare with human solutions. One of the first benchmarks to systematically evaluate autonomous ML agents on Kaggle competitions was MLAgentBench (Huang et al. 2023), which focused on a small set of tasks with simple baselines, measuring agents’ ability to improve upon them. Later, DSBench (Jing et al. 2024), expanded the scope but often relied on automated filtering, which excluded many complex or non-standard competitions. The most recent effort, MLE-Bench (Chan et al. 2024) stands out for its scale and diversity, present a more challenging and realistic testbed for multi-agent AutoML systems. However, MLE-Bench also faces notable limitations, including its large size (3.3 TB) and the fact that it constructs its test sets by partitioning the original training data and then compares the resulting scores to the private leaderboard positions that cause evaluation bias.
3 KompeteAI
The pipeline of KompeteAI is demonstrated in Figure 1. It is designed to ensure robust, leak-free data handling, efficient exploration of modeling ideas, and rapid iteration, all while maintaining high code quality and logical consistency. This section outlines the key stages and mechanisms that underpin the operation of our system. It consists of four main components: Pipeline Setup, The Ideation Process, Tree-Guided Exploration, which includes operations such as node addition and merging, and the Scoring Model, designed to accelerate overall pipeline evaluation.
In our framework, each node represents a segment of the final code, and each tree level corresponds to a distinct pipeline component: Exploratory Data Analysis (EDA), Feature Engineering (FE), or Model Training (MT). Connections between nodes indicate that they form parts of the same pipeline. The main stage unfolds in three phases: tree initialization, adding, and merging. Tree initialization constructs the primary tree structure. Subsequently, within the allotted time budget, the system alternates between the adding phase — where new ideas are injected into the most promising branches — and the merging phase, which combines two promising ideas into a single, more powerful solution. The EDA phase is represented by a single root node, which aggregates all insights derived from data exploration, such as visualizations, distribution analyses, and summary statistics. This EDA node is not static: as the search progresses, it can be dynamically enriched with additional analyses prompted by downstream discoveries. In contrast, nodes at the FE and MT levels represent concrete instantiations of their respective phases, with edges between nodes indicating their inclusion in the same candidate solution. While a node may have multiple children, it is restricted to a single parent.
3.1 Pipeline Setup
This phase sets up the core components required for the next stages. The dataset is ingested by the Reader Agent, which analyzes its structure, produces a detailed task specification, and initializes the data for the RAG based on this description. The Metric Agent constructs unit tests to support submission validation and defines the evaluation metric function. The Validator Agent partitions the data according to the task specification and applies appropriate preprocessing methods to ensure a valid and reliable evaluation protocol. The Baseliner Agent generates an initial solution, establishing a lower-bound reference for expected performance. Based on the baseline score, it also assesses the quality of the data split and, if necessary, can trigger the Validator Agent to re-partition the data using an alternative strategy.
3.2 The Ideation Process
This section focuses on the generation process for a single node within the pipeline, emphasizing localized decision-making rather than full-path construction. The procedure is structured as a modular interaction among four specialized agents. The Insighter Agent initiates the process by proposing candidate ideas. These are then evaluated by the Checker Agent for consistency and compatibility with the solution context. Once validated, the Coder Agent translates the idea into executable code. The output is again verified by the Checker Agent, this time to ensure implementation correctness. Finally, the Debugger Agent addresses potential runtime errors and integration issues, completing the node’s initialization cycle.
Figure 1: The KompeteAI AutoML pipeline consists of three main stages: Pipeline Setup, Tree Initialization, and the Main Loop. Pipeline Setup involves data ingestion, initial analysis, metric evaluation, data validation, and establishing performance benchmarks. Tree Initialization uses EDA to generate initial insights, forming a tree structure where Feature Engineering nodes are parents to Model Training nodes. The Main Loop iteratively refines the solution tree through adding new nodes, with their performance subsequently evaluated via code acceleration, debugging, and return speed to get a score; backpropagation for scoring FE ideas based on downstream model feedback; and merging strong partial solutions to recombine effective strategies
Insighter
The Insighter Agent generates ideas for downstream components, guided by factors such as EDA results and the current solution stage. To overcome LLM limitations in diversity and quality, it uses two key mechanisms.
  • Tree Memory: The agent employs a memory module over the entire ideation tree, analyzing previously generated nodes using diversity-driven strategies based on embedding similarity — such as selecting nodes closest or farthest from the parent, or sampling randomly. A top-$n$ subset is selected for inclusion in the context, and the memory is continuously updated as new nodes are added, enabling dynamic promotion of diversity.
  • Retrieval-Augmented Generation: The RAG mechanism retrieves strong ideas from Kaggle (top-$n$ similar tasks and winning solutions) and arXiv (top-$m$ papers via task-specific queries), forming a candidate pool. It then normalizes texts, extracts key ideas, and selects $k$ for context. The retrieval approach is adaptive, based on function calling that is triggered only when external knowledge is expected to contribute meaningfully, thereby minimizing computational overhead.
Checker
The Checker Agent is designed to validate the logical consistency of outputs produced by other agents. It uses unit tests, including schema validation and script execution, to assess correctness.
Coder
The Coder Agent implements the input idea based on several parameters, including a description of the input data, available computational resources, and the idea itself. It selects optimal hyperparameters and designs the architecture to ensure the code executes within the given time constraints while remaining as efficient as possible.
Debugger
The Debugger agent efficiently resolves issues with dependency installation, code generation, and submission formatting using a nested loop that iteratively debugs code within a set limit. To reduce runtime overhead common in existing systems, it accelerates debugging by minimizing time-sensitive parameters like training iterations, enabling fast error detection without full execution. Once debugging succeeds, the original configuration is restored. The agent also logs detailed metrics — such as error types, retries, and outcomes — and skips debugging steps for recurring errors, opting instead for direct code regeneration.
3.3 Tree-Guided Exploration
This section provides an explanation of the core principles underlying operations on the ideation tree.
Tree Initialization
The primary objective of the tree initialization stage is to generate an initial set of candidate pipelines, providing a diverse foundation from which further exploration can proceed. This stage seeds the search space with promising ideas and establishes the initial structure for subsequent optimization. After the initial tree is constructed, we perform a backpropagation step to propagate performance signals from the model training nodes (where direct evaluation scores are available) up through the tree, updating each node’s average score based on new information to improve their representativeness within the search tree.
Adding
The adding stage governs the structured expansion of the ideation tree $T_{t}=(V_{t},E_{t})$ by proposing new nodes at the Feature Engineering and Model Training levels. Before expansion, the agent examines the current tree state and may selectively trigger additional exploratory analysis via function calling, choosing specific data inspections that are most informative for subsequent node generation.
The entire process is conditioned on a global context representation $\mathbf{c}_{t}$, which captures the current pipeline state and external knowledge in a structured and denoised form. We define the context vector as:
$\mathbf{c}_{t}=\phi_{\mathrm{EDA}}(T_{t})+\phi_{\mathrm{reader}}(T_{t})+\phi_{\mathrm{ext}}(\texttt{QueryExternal}())$
where each mapping $\phi:\mathcal{X}\rightarrow\mathcal{C}$ transforms raw, unstructured, or noisy inputs into structured semantic representations in a common context space $\mathcal{C}$. Specifically:
  • $\phi_{\mathrm{EDA}}$ encodes statistical and structural information extracted from exploratory data analysis;
  • $\phi_{\mathrm{reader}}$ derives contextual insights from competition metadata and prior solutions;
  • $\phi_{\mathrm{ext}}$ incorporates knowledge retrieved from external sources (e.g., Kaggle, arXiv).
Based on the aggregated context $\mathbf{c}_{t}$, the agent performs a structured expansion of the ideation tree in three successive steps. First, a set of candidate FE nodes is sampled from the conditional distribution
$q_{\mathrm{FE}}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{V}_{\mathrm{FE}}),\quad q_{\mathrm{FE}}(\cdot\mid\mathbf{c}_{t})$
where $\mathcal{V}_{\mathrm{FE}}$ denotes the space of available FE transformations. Then, for each sampled FE node $v^{\mathrm{FE}}\in\mathcal{V}_{\mathrm{FE}}$, the agent generates a corresponding set of MT nodes by sampling from
$q_{\mathrm{MT}}:\mathcal{V}_{\mathrm{FE}}\rightarrow\mathcal{P}(\mathcal{V}_{\mathrm{MT}})$
which specifies a distribution over the MT configuration space $\mathcal{V}_{\mathrm{MT}}$. Finally, a subset of FE nodes from the current tree (including both existing and newly added ones) is selected based on their scores, reflecting their estimated utility. For each selected FE node, additional MT nodes are appended as children.
The full procedure is detailed in Appendix A1.
Merging
The merging stage enables the agent to consolidate multiple promising solutions at the Feature Engineering and Model Training levels, yielding stronger and more generalizable configurations.
The merging process unfolds as follows.
  1. 1.
    A set of FE node pairs $(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$ is sampled, excluding those present in the long-term memory buffer $\mathcal{M}_{\mathrm{long}}$. Each valid pair is merged into a new node
    $v_{ij}^{\mathrm{FE}}=\texttt{MergeFE}(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$
    which recombines structural and statistical traits of its parents.
  2. 2.
    For each merged node $v_{ij}^{\mathrm{FE}}$, a set of child MT nodes is generated from a conditional distribution. Then, for each parent FE node $v_{i}^{\mathrm{FE}}$ and $v_{j}^{\mathrm{FE}}$, the agent selects additional MT nodes from their respective subtrees. These are sampled stochastically with probabilities proportional to their scores:
    $u_{k}^{(i)}\sim\texttt{SampleTop}(v_{i}^{\mathrm{FE}})$
    The final child set of $v_{ij}^{\mathrm{FE}}$ combines the freshly generated MT nodes and the resampled top performers from its parents.
  3. 3.
    Additionally, a subset $\mathcal{V}_{\mathrm{FE}}^{\mathrm{merge}}\subseteq V_{t}^{\mathrm{FE}}$ is selected, and within each selected FE node, MT child pairs are merged using
    $u_{ij}^{\mathrm{MT}}=\texttt{MergeMT}(u_{i},u_{j})$
    to form stronger model configurations.
    To avoid redundant or destructive merges, the agent employs dual memory buffers. The short-term memory $\mathcal{M}_{\mathrm{short}}$ temporarily stores failed merge attempts, while the long-term memory $\mathcal{M}_{\mathrm{long}}$ permanently excludes repeatedly failing node pairs. Specifically, a pair that fails to produce a beneficial merge $\theta_{\text{fail}}$ times is promoted from $\mathcal{M}_{\mathrm{short}}$ to $\mathcal{M}_{\mathrm{long}}$, ensuring efficient resource allocation and adaptive learning.
Merging procedure detailed in Appendix A2.
3.4 Scoring Model
The scoring model is a crucial component applied at the model training nodes of our pipeline generation tree. Its primary purpose is to accelerate the overall pipeline evaluation process, thereby enabling the exploration of a greater number and diversity of ideas within the solution space. By providing rapid performance estimates for candidate models, the scoring model allows us to prioritize the most promising approaches without the need for full-scale training, which can be prohibitively time-consuming for complex models.
The core principle behind the scoring model is to predict the final performance of a candidate model based on how similar models have performed on the same dataset. To achieve this, we generate detailed descriptions for several anchor ideas as well as for the candidate idea to be scored. This step is essential, as even minor code differences — such as the number of training epochs — can lead to significant variations in model performance. These descriptions are then inserted into a prompt to LLM with dataset description and current model training specification.
To construct effective anchor ideas for the scoring model, we first train several model architectures on a single feature engineering idea to observe how different architectures behave on the dataset. Subsequently, we select one model architecture and train it across all feature engineering nodes, allowing us to assess the impact of each feature engineering strategy on the final model score. This systematic approach to few-shot example selection ensures that the scoring model receives diverse and informative context, further enhancing its predictive accuracy.
4 Kompete-bench
4.1 Motivation
Despite growing interest in multi-agent AutoML, empirical evaluation remains limited by the lack of accessible and reliable benchmarks. Currently, the only public option is MLE-Bench — a collection of 75 Kaggle competitions that offer a rich mix of real-world ML challenges, clear evaluation metrics, and a competitive setup that closely mirrors practical deployments.
While MLE-Bench is a significant step forward, it has two key limitations that hinder its practical utility and reproducibility. First, it is prohibitively large: even the ”Lite” version with just 22 competitions takes up 158GB and requires substantial compute to process. Second, MLE-Bench constructs test sets by partitioning the original training data and compares results to private leaderboard positions. This introduces significant evaluation bias, as it does not reflect actual test set performance. We show that scores on MLE-Bench’s constructed test sets can diverge significantly from real leaderboard rankings, leading to misleading conclusions and undermining its reliability as a proxy for real-world outcomes.
4.2 Methodology
We propose a concise two-part benchmark for evaluating multi‐agent AutoML systems, balancing historical, contemporary, and future‐oriented challenges under standardized computational constraints.
  • Selection of Established Competitions. We curate 15 competitions from the ’lite’ MLEBench collection that still accept late submissions on Kaggle and whose individual dataset sizes do not exceed 1 GB. The aggregate volume of this subset is 5.3 GB, providing a stable foundation for baseline comparisons.
  • Incorporation of New Competitions. To better reflect the evolving landscape of AutoML tasks, we include 11 additional competitions from 2024 and 2025 years. These datasets, totaling 4.9 GB, were selected primarily to ensure fair comparison with human performance. This is because, given the recency of these competitions, both human participants and current models have access to almost the same tools and libraries, minimizing discrepancies due to technological advancements.
5 Experiment
5.1 Experiment Setup
Baselines
To provide a comprehensive evaluation on MLE-Bench, we compared our system against the top-ranked methods on the leaderboard: AIDE, RD-agent, and Ml-Master. Due to the high cost of running all baselines, we report their leaderboard metrics as provided directly by MLE-Bench. For evaluation on Kompete-bench we re-ran only AIDE and RD-agent, as both are open-source, while the implementation of Ml-Master was not publicly available at the time of our study and was therefore excluded. All methods were tested using the same LLM backend — gemini-2.5-flash (Comanici et al. 2025), which also powers our system. Additionally, to fairly assess the alignment between medal distributions on MLE-Bench and real Kaggle leaderboards, we also evaluated all methods using o1-preview (Jaech et al. 2024), the model on which most MLE-Bench submissions achieved their highest scores.
Environment
To evaluate our system on MLE-Bench, we set a 6-hour runtime limit using 12 vCPUs, 64 GB of RAM, and a single NVIDIA A100 GPU (40 GB). The same hardware setup was used for all systems on Kompete-bench, where we applied a stricter 6-hour time limit. Each configuration was executed three times, and results were averaged to reduce variance. To ensure fairness and prevent systems from exploiting test-time information, we additionally RAG components from accessing any information about the competition date on which the agent was evaluated.
Benchmarks
We evaluate KompeteAI on the Lite subset of MLE-Bench, a computationally feasible proxy for the full benchmark. Although not identical, MLE-Bench reports high alignment in system rankings between Lite and full versions, making Lite a practical basis for comparison. For medal-based evaluation against real Kaggle leaderboards, we include only Lite competitions that still accept submissions. Additionally, we assess performance on our custom benchmark — Kompete-bench, structured following the same principles described in our methodology section, by separating the data into two categories: MLE-subset Lite and recent and a collection of newly curated competitions.
Evaluation Metrics
On MLE-Bench, we adopt the official leaderboard metric — the percentage of submissions that receive a medal — which aligns with standard Kaggle competition criteria. For Kompete-bench, we additionally report the percent humans beaten metric, which measures the percentage of human participants outperformed by the agent on the corresponding Kaggle leaderboard. This metric offers a more fine-grained and interpretable evaluation signal compared to medal thresholds, allowing us to distinguish between systems that may not earn medals but differ meaningfully in their relative competitiveness against human participants. To assess pipeline efficiency, we report the number of complete pipeline passes, where a single pass corresponds to one adding and one merging iteration performed by the AutoML system.
5.2 Results
Evaluation on MLE-Bench
We begin by evaluating our system on the widely adopted MLE-Bench benchmark (Lite subset). As presented in Table 2, KompeteAI attains the highest overall performance, outperforming existing state-of-the-art systems by an average margin of 3 percentage points.
Name MLE-Bench medals %
AIDE (o1-preview) 34.3 ± 2.4
RD-agent (o1-preview) 48.18 ± 2.49
ML-Master (deepseek-r1) 48.5 ± 1.5
KompeteAI (gemini-2.5-flash) 51.5 ± 1.5
Table 2: Comparison of agents on MLE-Bench ’Lite’ subset. For AIDE and RD-agent, we use results reported by MLE-Bench. KompeteAI was run 3 times with different seeds; results are reported as mean ± SEM. Each run was given 6 hours.
Kaggle-Based Validation of Robustness
To validate robustness beyond proxy benchmarks, we tested all agents on the Lite subset using official Kaggle submissions (Table 3). The analysis reveals a substantial gap between MLE-Bench evaluations and actual leaderboard outcomes. In particular, both AIDE and RD-agent exhibit inflated medal rates on MLE-Bench compared to their real Kaggle performance. This discrepancy highlights the limitations of relying solely on MLE-Bench for comparison, especially in competitive human-level settings.
Name MLE-Bench LB Real LB Contemporary part
AIDE (o1-preview) 20 7 0
RD-agent (o1-preview) 27 13 0
AIDE (gemini-2.5-flash) 13 7 0
RD-agent (gemini-2.5-flash) 13 7 0
Table 3: Medal rate gap between MLE-Bench leaderboard and leaderboard of real Kaggle competition. For this experiment we ran each system once and used the rules from Kaggle to determine the medal.
Evaluation on Kompete-Bench
To address these limitations, Kompete-Bench evaluates performance using real Kaggle leaderboards, covering recent competitions with high prize pools and strong engagement. On this Contemporary subset, traditional medal-based evaluation fails, as all agents achieve zero medal rates. We therefore propose a finer-grained metric: percent humans beaten. As shown in Figure 2, KompeteAI demonstrates state-of-the-art results, surpassing human performance in 11.2% of cases — significantly ahead of both RD-agent and AIDE. Nevertheless, the agent remains far behind top leaderboard teams, reflecting how contemporary competitions demand capabilities beyond local modeling — including large-scale feature engineering, synthetic data generation, and use of external datasets. Human participants leverage broader tools and resources, underscoring the competitiveness of these challenges and the remaining headroom for automated agents.
Impact of the Acceleration Methods
Finally, we assess the contribution of our acceleration framework. Table 4 reports the number of iterations within a fixed budget. By integrating predictive scoring and an accelerated debugging loop, our pipeline executes over 6.9× more iterations than a baseline without accelerations, enabling stronger submissions under identical constraints. Moreover, as shown in Figure 2, KompeteAI without accelerations exhibits a marked performance drop, underscoring its ability to refine solutions during test-time optimization. While we restrict acceleration measurements to KompeteAI due to tight pipeline integration, the proposed paradigms are general and may be adapted to other architectures with minimal conceptual changes.
Figure 2: Comparison of our pipeline with AIDE and RD-agent on Contemporary and MLE-Bench parts of Kompete-bench. All systems use gemini-2.5-flash as the underlying LLM. Each was run 3 times with different seeds, results are averaged, and each run was limited to 6 hours.
System Configuration Number of Iterations
[…]
with all accelerations 12.5 ± 2.1
Table 4: Impact of acceleration techniques on the number of completed iterations within a fixed time budget. Mean ± 95% CI estimated via Student’s t-distribution.
5.3 Ablation Study
As shown in Figure 2, all major components of KompeteAI are crucial, with their impact particularly pronounced on the Contemporary subset.
[…]
6 Conclusions
We introduced KompeteAI, a new multi-agent AutoML framework, and demonstrated its strong performance across diverse and challenging tasks. Our work opens several promising directions for future research.
First, while our acceleration paradigm significantly reduces computation time, it relies on a scoring model whose accuracy may degrade over longer runs, potentially leading to cumulative errors. Improving this component — e.g., through adaptive retraining or uncertainty-aware corrections — could enhance long-term robustness.
Second, we see strong potential in deepening the integration of LLM-based reasoning with algorithmic search. By embedding structured computation within a language-guided planning process, agents can explore solution spaces more effectively. Extending this to coordinated multi-agent fine-tuning — with shared representations and task-aware interaction — could further strengthen system-level coherence and adaptability.
Finally, systems like KompeteAI may prove valuable beyond competitive AutoML settings. While platforms like Kaggle offer practical, open-ended challenges, they only partially reflect the demands of real-world scientific discovery. As autonomous agents begin contributing to hypothesis generation, experimental design, and even paper drafting, refining these tools for real research workflows becomes an exciting and urgent frontier.
Acknowledgements
The study research has been supported by the Ministry of Economic Development of the Russian Federation (agreement No. 139-10-2025-034 dd. 19.06.2025, IGK  [^00] C313925P4D0002).
Appendix A Algorithms
A.1 Adding
This subsection formalizes the adding stage, as defined in Algorithm 1. We define the ideation tree at iteration $t$ as $T_{t}=(V_{t},E_{t})$, and the global context vector as $\mathbf{c}_{t}\in\mathcal{C}$, where $\mathcal{C}$ denotes the space of structured semantic representations.
The agent samples new Feature Engineering (FE) and Model Training (MT) nodes from conditional distributions:
$q_{\mathrm{FE}}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{V}_{\mathrm{FE}}),\quad q_{\mathrm{MT}}:\mathcal{V}_{\mathrm{FE}}\rightarrow\mathcal{P}(\mathcal{V}_{\mathrm{MT}})$
where $\mathcal{V}_{\mathrm{FE}}$ and $\mathcal{V}_{\mathrm{MT}}$ are the respective candidate spaces for FE and MT nodes, and $\mathcal{P}(\cdot)$ denotes the space of probability distributions over a discrete set.
We also define a scoring function $a:\mathcal{V}_{\mathrm{MT}}\rightarrow\mathbb{R}$, and a softmax-based selection mechanism used to stochastically expand the MT subtrees. The resulting tree $T_{t+1}$ is obtained by applying the additions and structural updates defined by the algorithm.
Algorithm 1 Adding Stage at Iteration $t$
Input: Ideation tree $T_{t}=(V_{t},E_{t})$; context vector $\mathbf{c}_{t}$
Parameters: $N$ (number of FE nodes), $M$ (number of MT nodes per parent)
Output: Updated tree $T_{t+1}$
[…]
4:for each $v_{j}^{\mathrm{FE}}$ do
5:  $\{u_{j,i}^{\mathrm{MT}}\}_{i=1}^{M}\sim q_{\mathrm{MT}}(\cdot\mid v_{j}^{\mathrm{FE}})$
6:  Assign scores $\{a_{j,i}\}_{i=1}^{M}$
7:end for
[…]
9: Transform scores:
$s_{j,i}=\begin{cases}a_{j,i},&\text{if higher is better}\\ -a_{j,i},&\text{otherwise}\end{cases}$
10: Compute probabilities:
$p_{j,i}=Softmax(s_{j,i})$
11: Sample subset: $\mathcal{S}\sim\texttt{Sample}(\{u_{j,i}^{\mathrm{MT}}\},\{p_{j,i}\})$
12:for each $u\in\mathcal{S}$ do
13:  $\{w_{k}^{\mathrm{MT}}\}_{k=1}^{M}\sim q_{\mathrm{MT}}(\cdot\mid u)$
14:end for
[…]
16:return $T_{t+1}$
A.2 Merging
This subsection defines the merging stage and its associated memory mechanism. Let
$\mathcal{M}_{\mathrm{short}},\mathcal{M}_{\mathrm{long}}\subseteq\mathcal{P}(V_{t}^{\mathrm{FE}})$
denote short- and long-term memory buffers tracking merge success.
Algorithm 2 Merging Stage at Iteration $t$
Input: Ideation tree $T_{t}=(V_{t},E_{t})$; memories $\mathcal{M}_{\mathrm{short}},\mathcal{M}_{\mathrm{long}}$
Parameters: $N$ (number of FE merge candidates), $M$ (MT children per merged FE), $\theta_{\mathrm{fail}}$ (failure threshold)
Output: Updated tree $T_{t+1}$, updated memories
[…]
2:for each $(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$ do
3:  $v_{ij}^{\mathrm{FE}}\leftarrow\texttt{MergeFE}(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$
4:  $\{u_{ij,k}^{\mathrm{MT}}\}_{k=1}^{M}\sim q_{\mathrm{MT}}(\cdot\mid v_{ij}^{\mathrm{FE}})$
5:  $\mathcal{S}_{i},\mathcal{S}_{j}\sim\texttt{SampleTop}(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$
6:  Attach $\{u_{ij,k}^{\mathrm{MT}}\}\cup\mathcal{S}_{i}\cup\mathcal{S}_{j}$ to $v_{ij}^{\mathrm{FE}}$
7:  $\Delta_{ij}\leftarrow\texttt{Evaluate}(v_{ij}^{\mathrm{FE}})$
8:  if $\sum_{r=1}^{R}\mathbb{I}\big{[}\texttt{IsFailure}(\Delta_{ij}^{(r)})\big{]}\geqslant\theta_{\mathrm{fail}}$ then
9:   $(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})\in\mathcal{M}_{\mathrm{long}}$
10:  else
[…]
12:    $(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})\in\mathcal{M}_{\mathrm{short}}$
13:   else
14:    $(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})\in\mathcal{M}_{\mathrm{long}}$
15:   end if
16:  end if
17:end for
[…]
19:for each $v^{\mathrm{FE}}\in\mathcal{V}_{\mathrm{FE}}^{\mathrm{merge}}$ do
20:  $(u_{p}^{\mathrm{MT}},u_{q}^{\mathrm{MT}})\sim q_{\mathrm{pair}}(\cdot\mid\texttt{Children}(v^{\mathrm{FE}}))$
21:  $u_{pq}^{\mathrm{MT}}\leftarrow\texttt{MergeMT}(u_{p}^{\mathrm{MT}},u_{q}^{\mathrm{MT}})$
22:  Attach $u_{pq}^{\mathrm{MT}}$ to $v^{\mathrm{FE}}$
23:end for
[…]
25:return $T_{t+1},\mathcal{M}_{\mathrm{short}},\mathcal{M}_{\mathrm{long}}$
Appendix B Kompete-bench
Figure A3: End dates for competitions presented in Kompete-Bench
Full descriptions of the competitions in Kompete-bench, including the name, number of participants, metrics, and medal thresholds, are listed in Table 5. The benchmark comprises 26 Kaggle competitions, totaling 10.2 GB in size, and is divided into two distinct parts. The first part includes competitions from MLE-Bench that remain open for submissions on Kaggle and are each under 1 GB. These span from 2014 to 2017 and primarily feature straightforward tasks, where strong leaderboard positions can be achieved without complex modeling or novel ideas — sometimes even leveraging tools and libraries unavailable at the time of the original competition.
The second part consists of more recent competitions from 2024–2025, some of which are still ongoing. Achieving a medal in these requires creative approaches or training a large number of models, reflecting the increased difficulty and evolving standards of modern machine learning challenges.
The distribution of competition end dates is illustrated in Figure A3, clearly showing the temporal separation between the two parts of the benchmark.
For all competitions, we report the percentage of participants outperformed (“percent users beaten”) as the primary metric. This choice is motivated by the observation that current AutoML systems are unable to achieve medal positions in the latest competitions, yet tracking progress on these challenging tasks remains crucial. The ”percent users beaten” metric is computed using the actual private Kaggle leaderboard, averaged over three independent runs. If a system fails to generate a valid submission, a score of 0% is assigned for that run. For ongoing competitions, evaluation is performed simultaneously on the public leaderboard.
Name Number of participants Metric Bronze Silver Gold Part aerial-cactus-identification 1221 ROC-AUC ↑ 1 1 1 MLE-bench (Lite) denoising-dirty-documents 162 RMSE ↓ 0.04517 0.02609 0.01794 MLE-bench (Lite) dog-breed-identification 1281 log loss ↓ 0.04598 0.00539 0.0005 MLE-bench (Lite) dogs-vs-cats-redux-kernels-edition 1315 log loss ↓ 0.06127 0.05038 0.03882 MLE-bench (Lite) jigsaw-toxic-comment-classification-challenge 4539 mean col-wise ROC AUC ↑ 0.98639 0.98668 0.98740 MLE-bench (Lite) leaf-classification 1596 log loss ↓ 0.01526 0.00791 0.00000 MLE-bench (Lite) mlsp-2013-birds 81 ROC-AUC ↑ 0.87372 0.90038 0.93527 MLE-bench (Lite) nomad2018-predict-transparent-conductors 879 RMSLE ↓ 0.06582 0.06229 0.05589 MLE-bench (Lite) plant-pathology-2020-fgvc7 1318 ROC-AUC ↑ 0.97361 0.97465 0.97836 MLE-bench (Lite) random-acts-of-pizza 462 ROC-AUC ↑ 0.6921 0.76482 0.97908 MLE-bench (Lite) spooky-author-identification 1242 log loss ↓ 0.29381 0.26996 0.16506 MLE-bench (Lite) tabular-playground-series-dec-2021 1189 ROC-AUC ↑ 0.95658 0.95658 0.9566 MLE-bench (Lite) tabular-playground-series-may-2022 1152 ROC-AUC ↑ 0.99818 0.99822 0.99823 MLE-bench (Lite) text-normalization-challenge-english-language 261 accuracy ↑ 0.99038 0.99135 0.99724 MLE-bench (Lite) text-normalization-challenge-russian-language 163 accuracy ↑ 0.97592 0.98232 0.99012 MLE-bench (Lite) eedi-mining-misconceptions-in-mathematics 1449 MAP@25 ↑ 0.46090 0.49136 0.56429 Contemporary learning-agency-lab-automated-essay-scoring-2 2708 quadratic weighted kappa ↑ 0.83471 0.83518 0.83583 Contemporary lmsys-chatbot-arena 1688 log loss ↓ 1.00472 0.99410 0.98392 Contemporary pii-detection-removal-from-educational-data 2049 efficiency score ↑ 0.95714 0.95883 0.96615 Contemporary um-game-playing-strength-of-mcts-variants 1610 RMSE ↓ 0.43050 0.42973 0.42591 Contemporary llm-prompt-recovery 2176 Sharpened Cosine Similarity ↑ 0.6375 0.6513 0.6848 Contemporary equity-post-HCT-survival-predictions 3327 C-index ↑ 0.69288 0.69320 0.69500 Contemporary cmi-detect-behavior-with-sensor-data 2156 F1 ↑ 0.84 0.84 0.86 Contemporary make-data-count-finding-data-references 833 F1 ↑ 0.548 0.564 0.620 Contemporary neurips-open-polymer-prediction-2025 1539 wMAE ↓ 0.057 0.041 0.032 Contemporary wsdm-cup-multilingual-chatbot-arena 890 categorization accuracy ↑ 0.696381 0.702772 0.711412 Contemporary
Table 5: Summary of tasks and metrics included in Kompete-bench: the table lists competition names, number of participants, evaluation metrics, threshold values for medals and the part of the benchmark to which this competition belongs
Appendix C Setup details
We used the hyperparameter settings summarized in Table 8. For AIDE and the RD-agent, we retained their default configurations as specified in the original implementations except for the time limit. For KompeteAI, we empirically selected a set of hyperparameters that strike a balance between the quality of component exploration and the computational time allocated to each. This tuning was guided by the need to ensure efficient coverage of critical components while maintaining tractable execution time. Below are the hyperparameters and their descriptions for each system.
KompeteAI
The list of hyperparameters used when running KompeteAI on benchmarks is given in the Table 6.The time_run_minutes parameter sets the maximum runtime for the entire multi-agent system in minutes, after which the process will terminate. The runtime_error_time defines the time limit (in minutes) after which a generated code is will be stopped. The subset_size_in_percent specifies the percentage of the dataset to be used for quick validation. The validator_size_threshold sets a threshold for the dataset size; if the data exceeds this value, a subset is used for training and validation. The number_of_ideas_eda determines how many exploratory data analysis (EDA) ideas generates per iteration. Similarly, number_of_ideas_data and number_of_ideas_modelling control the number of ideas related to feature engineering and model training. The max_add_idea parameter limits how many new ideas can be added to the idea pool in a single adding iteration. The number_of_selected_node specifies how many nodes are selected for adding expansion at each step. The number_of_iterations_parents sets how many iterations parent nodes participate in generating new ideas, while number_of_selected_node_merging determines how many nodes are chosen for merging at each iteration. The number_of_ideas_min and number_of_ideas_max define the minimum and maximum number of ideas, which are used as achor ideas for scoring model. The retrieve_n_papers and retrieve_n_competitions parameters control how many papers retrieves from arXiv papers and Kaggle solutions. The number_rag_ideas sets how many ideas are generated using RAG. The memory_size parameter determines how many recent ideas, solutions, or states each agent remembers for learning and decision-making. Alternatively, if memory_size is set to nearest_nodes, the agent’s memory consists of the most similar nodes or ideas, rather than a fixed number, allowing for more contextually relevant recall.
Hyperparam name Value time_run_minutes 360 runtime_error_time 30 subset_size_in_percent 10 validator_size_threshold $10^{4}$ number_of_ideas_eda 5 number_of_ideas_data 2 number_of_ideas_modelling 2 max_add_idea 2 number_of_selected_node 2 number_of_iterations_parents 2 number_of_selected_node_merging 2 number_of_iterations_children 3 number_of_ideas_min 2 number_of_ideas_max 5 retrieve_n_papers 3 retrieve_n_competitions 3 number_rag_ideas 5
Table 6: Hyperparameters used for running KompeteAI
RD-Agent
The list of hyperparameters used when running RD-agent on benchmarks is given in the Table 7. The debug_timeout parameter sets the maximum time, in seconds, that is allowed to debug one generated code. The full_timeout parameter defines the overall time limit for system. The if_action_choosing_based_on_UCB flag determines whether the agents select their actions using the Upper Confidence Bound (UCB) strategy. The enable_knowledge_base flag indicates whether a shared knowledge base is enabled for the agents. The loop_n parameter specifies the number of main iterations the system will execute, set to 2000 to avoid early stopping not by the time limit. Finally, the max_trace_num parameter limits the number of traces that can be created during system execution.
Hyperparam name Value debug_timeout 3600 full_timeout 21600 if_action_choosing_based_on_UCB False enable_knowledge_base False loop_n 2000 max_trace_num 3
Table 7: Hyperparameters used for running RD-agent
AIDE
The list of hyperparameters used when running RD-agent on benchmarks is given in the Table 8. The steps parameter defines the maximum number of steps the entire system can perform during run. The max_debug_depth, specifies the maximum depth for recursive code debugging. The debug_prob parameter suggests that debugging is enabled for every generated code, meaning that all relevant information will be recorded without any sampling. Finally, the time_limit parameter represents the maximum allowed wall-clock time (in seconds) for experiment.
Hyperparam name Value steps 2000 max_debug_depth 20 debug_prob 1 time_limit 21600
Table 8: Hyperparameters used for running AIDE
Appendix D Metrics
Competition RD-agent AIDE KompeteAI aerial-cactus-identification 58 5 79 denoising-dirty-documents 3 None 12 dog-breed-identification 41 62 58 dogs-vs-cats-redux-kernels-edition 86 6 91 jigsaw-toxic-comment-classification-challenge 42 26 88 leaf-classification 38 24 46 mlsp-2013-birds None 0 0 nomad2018-predict-transparent-conductors 78 33 22 plant-pathology-2020-fgvc7 12 62 68 random-acts-of-pizza 58 47 67 spooky-author-identification 71 47 59 tabular-playground-series-dec-2021 17 22 25 tabular-playground-series-may-2022 16 32 36 text-normalization-challenge-english-language 3 9 27 text-normalization-challenge-russian-language 17 14 18
Table 9: Full table by ”percent humans beaten” for AutoML systems given in the article at each competition in MLE-Bench part of Kompete-bench
Competition RD-agent AIDE KompeteAI cmi-detect-behavior-with-sensor-data None None None eedi-mining-misconceptions-in-mathematics 22 None 30 equity-post-HCT-survival-predictions None None None learning-agency-lab-automated-essay-scoring-2 12 19 13 llm-prompt-recovery 12 9 None lmsys-chatbot-arena None None None make-data-count-finding-data-references 11 None 18 neurips-open-polymer-prediction-2025 13 None 15 pii-detection-removal-from-educational-data None None 29 um-game-playing-strength-of-mcts-variants None None None wsdm-cup-multilingual-chatbot-arena 14 None 19
Table 10: Full table by ”percent humans beaten” for AutoML systems given in the article at each competition in Contemporary part of Kompete-bench.
Results of AutoML system on Kompete-bench are presented in the Table 9 and Table 10. We report results using the ”percent humans beaten” metric, which reflects the percentage of Kaggle leaderboard participants outperformed by each system. For every competition, each system was evaluated over three independent runs; the final score is the arithmetic mean of these runs, rounded to the nearest integer. If a system failed to generate any valid solution across all three attempts, its score is reported as None (equivalent to 0 when averaging across competitions). For individual runs where no valid submission was produced, a score of 0 was assigned. Notably, the primary factor influencing overall performance was the proportion of valid submissions: systems that consistently generated correct code achieved substantially higher scores.

### v2

@@
 We further expand the hypothesis space by integrating Retrieval-Augmented Generation (RAG), sourcing ideas from Kaggle notebooks and arXiv papers to incorporate real-world strategies.
-KompeteAI also addresses the execution bottleneck via a predictive scoring model and an accelerated debugging method, assessing solution potential using early stage metrics to avoid costly full-code execution.
+KompeteAI also addresses the execution bottleneck via a predictive scoring model and an accelerated debugging method, assessing solution potential using early-stage metrics to avoid costly full-code execution.
 This approach accelerates pipeline evaluation 6.9 times.
@@
 Recent research has shifted towards the use of Large Language Models (LLM) as the core reasoning engine for AutoML frameworks, enabling the autonomous generation and testing of end-to-end pipelines that adapt to specific tasks (Li et al.
-2024; Jiang et al.
-2025; Chi et al.
-2024; Liu et al.
-2025; Trirat, Jeong, and Hwang 2024; Grosnit et al.
-2024; Yang et al.
-2025).
+(2024); Jiang et al.
+(2025); Chi et al.
+(2024); Liu et al.
+(2025); Trirat et al.
+(2024); Grosnit et al.
+(2024); Yang et al.
+(2025); Nam et al.
+(2025)).
 However, these approaches have critical limitations.
 Initial ”one-shot” generation can yield diverse ideas but lacks iterative refinement, so a single flawed component - like suboptimal feature engineering — can undermine the entire pipeline without a way to correct it.
-More adaptive frameworks based on Monte Carlo Tree Search (MCTS) address this by exploring a tree of potential solutions, but often use constrained exploration and struggle to recombine promising ideas from different branches.
-Even recent frameworks that combine exploratory search with LLM-based reasoning (Liu et al.
-2025), despite achieving state-of-the-art results on MLE-Bench (Chan et al.
-2024), are limited by architectural constraints.
-They may not systematically preserve or merge valuable insights from distant high-performing branches.
+More adaptive frameworks based on Monte Carlo Tree Search (MCTS) address this by exploring a tree of potential solutions but often use constrained exploration and struggle to recombine promising ideas from different branches.
+Even recent frameworks that combine exploratory search with LLM-based reasoning Liu et al.
+(2025), despite achieving state-of-the-art results on MLE-Bench Chan et al.
+(2024), are limited by architectural constraints.
+They may not systematically preserve or merge valuable insights from distant high-performing branches during the exploration process.
 As a result, valuable solution components can be prematurely discarded.
-Even an ideal internal recombination can only reshuffles known ideas.
-Retrieval‑Augmented Generation (RAG) breaks this limit by injecting external, domain-specific knowledge — enabling the agent to explore solutions beyond its pretrained hypothesis space.
-This potential remains largely untapped in current implementations.
-Most systems apply RAG only during early stages; as the problem shifts — from statistical feature engineering to physics-based simulation or chemistry-driven descriptors — the agent cannot retrieve fresh, relevant knowledge, leading to knowledge decay as the pipeline evolves.
-Beyond search and knowledge limitations, all frameworks suffer from a severe execution bottleneck.
-Validating a single solution requires full code execution, which can take hours for complex models.
-This issue becomes even more problematic during debugging, as errors late in the pipeline force a complete retraining for every fix.
-As a result, the slow feedback loop discourages major changes and presents a serious scalability challenge.
+Even perfect internal recombination only reshuffles known ideas.
+Retrieval-Augmented Generation (RAG) overcomes this by injecting external, domain-specific knowledge, enabling exploration beyond the pretrained hypothesis space.
+Yet current systems rarely exploit this potential: most apply RAG only in the early stages Trirat et al.
+(2024); Yang et al.
+(2025), causing retrieval to fail at providing fresh knowledge and leading to knowledge decay across the pipeline.
+All frameworks also face a severe execution bottleneck.
+Validating a single solution requires full code execution, often taking hours.
+Debugging worsens this: late errors force complete retraining, slowing feedback and discouraging major changes.
+This creates a serious scalability challenge.
 In this work, we introduce KompeteAI, an autonomous multi-agent framework for structured, multistage pipeline generation.
-The key innovations compared to prior approaches is presented in Table 1.
+The key innovations compared to prior approaches are presented in Table 1.
 To efficiently explore and exploit the search space, we employ two core operators.
@@
 We address execution bottlenecks through a predictive scoring model that prunes weak solutions early, alongside an accelerated debugging paradigm using simplified code and smaller data samples, dramatically shortening the feedback loop.
-Our experiments show that KompeteAI sets a new state-of-the-art on MLE-Bench, outperforming prior methods by an average of 3%, while accelerating the average test-time performance by a factor of 6.9.
+These innovations accelerate the average test-time performance by a factor of 6.9.
+Our experiments show that KompeteAI sets the new state-of-the-art on MLE-Bench, outperforming prior methods by an average of 3%.
 Finally, the only publicly available competitive benchmark MLE-Bench suffers from two key limitations.
@@
 Our primary contributions:
-  • Stage-Decomposed Multi-Agent Architecture: A state-of-the-art framework that partitions the ML workflow into discrete stages, enabling agents to specialize in focused tasks, dynamically integrate external knowledge sources to enhance exploration diversity, and systematically recombine optimal partial solutions through novel addition and merging operators.
+  • Stage-Decomposed Multi-Agent Architecture: A state-of-the-art framework that partitions the ML workflow into discrete stages, enabling agents to specialize in focused tasks, dynamically integrate external knowledge sources to enhance exploration diversity, and systematically recombine optimal partial solutions through novel adding and merging operators.
   • An Accelerated Evaluation and Debugging Paradigm: A two-part solution to the execution bottleneck, combining a predictive scoring model and a rapid debugging framework to drastically reduce validation time.
@@
 2 Related Work
-Classic AutoML
-Classic AutoML frameworks - TPOT, AutoGluon, AutoKeras, LightAutoML, and others (Olson and Moore 2016; Erickson et al.
-2020; Jin, Song, and Hu 2019; Vakhrushev et al.
-2021; Feurer et al.
-2022; LeDell and Poirier 2020; Thornton et al.
-2013) — automate data preprocessing, model selection, and hyperparameter tuning via heuristic search, ensembling, and Bayesian optimization.
+Classic AutoML.
+Classic AutoML frameworks - TPOT, AutoGluon, AutoKeras, LightAutoML, and others Olson & Moore (2016); Erickson et al.
+(2020); Jin et al.
+(2019); Vakhrushev et al.
+(2021); Feurer et al.
+(2022); LeDell & Poirier (2020); Thornton et al.
+(2013) — automate data preprocessing, model selection, and hyperparameter tuning via heuristic search, ensembling, and Bayesian optimization.
 Despite their effectiveness, these systems operate within static search spaces, require manual data preparation and adaptation for each new task, and lack the dynamic coordination and continual learning capabilities inherent to multi‑agent AutoML architectures.
-AutoML based LLM
-LLM-based AutoML systems have progressed rapidly, offering increasingly autonomous capabilities through dynamic coordination and iterative planning (Li et al.
-2024; Chi et al.
-2024; Jiang et al.
-2025; Trirat, Jeong, and Hwang 2024; Yang et al.
-2025; Liu et al.
-2025).
-To analyze and compare these systems meaningfully, we focus on four key aspects that consistently shape their performance and flexibility: exploration strategy, RAG, merge methods, and debugging techniques.
-These components address critical challenges related to how systems explore the design space, incorporate external knowledge, manage idea diversity, and implement solutions both efficiently and reliably.
-Notably, RD-Agent’s merge approach amounts to simple, uncontrolled recombination driven by LLM rather than a structured merge algorithm, which results in incoherent or suboptimal pipeline integrations.
-A summary of how existing systems approach these dimensions is provided in Table 1.
-Table 1: Comparison of LLM‑based AutoML systems.
-Phase‑wise chain: sequential, stage‑by‑stage pipeline construction; without context: searches that ignore information from parallel or prior branches; Multi‑stage expansion: KompeteAI’s staged node growth at every pipeline level; R&D‑phase: retrieval only during initial planning, versus Dynamic RAG: on‑the‑fly literature/web queries; Incremental debugging: iterative draft–debug–refine cycles, versus Multistage debugging: KompeteAI’s multi‑agent, stage‑based error checking; Simple recombination: implicit LLM‑driven fusion of ideas, versus Controlled merger: KompeteAI’s explicit algorithmic branch combination.
-Method Exploration Strategy RAG system Debugging method Merge Method
-AutoKaggle Phase-wise chain Three-stage phase & Unit tests
-SELA MCTS (UCT) without context
-AIDE MCTS (UCT) without context Incremental
+LLM-based AutoML.
+LLM-based AutoML systems have progressed rapidly, offering increasingly autonomous capabilities through dynamic coordination and iterative planning (AutoKaggle Li et al.
+(2024), SELA Chi et al.
+(2024), AIDE Jiang et al.
+(2025), AutoML-Agent Trirat et al.
+(2024), RD-agentYang et al.
+(2025), ML-Master Liu et al.
+(2025), MLE-STAR Nam et al.
+(2025)).
+To analyze and compare these systems meaningfully, we focus on four key aspects.
+The results are presented in Table 1.
+Notably, RD-Agent’s merging relies on uncontrolled LLM-driven recombination, often producing incoherent or suboptimal integrations, while MLE-STAR’s single, terminal merging step simply consolidates ideas without meaningfully improving solution exploration.
+Exploration RAG Debuging Merging
+SELA MCTS (UCT) w/o context
+AIDE MCTS (UCT) w/o context Incremental
 AutoML-Agent Retrieval-augmented planning R&D-phase Incremental
-R&D-Agent Multi-trace exploration R&D-phase Incremental Simple recombination
-ML-Master Controlled MCTS with memory Incremental
-KompeteAI Multi-stage expansion Dynamic Multistage Controlled merger
-Scoring Model
-Our approach to model scoring is inspired by similar performance prediction methods developed in Neural Architecture Search (Elsken, Metzen, and Hutter 2019).
-In NAS, a common approach is to use weight-sharing supernets, where multiple architectures are jointly trained by sharing parameters with a large model, and performance is estimated by evaluating sampled architectures on a validation set (Jawahar et al.
-2023a).
-This method, however, faces challenges such as weight co-adaptation (Bender et al.
-2018), capacity bottlenecks (Jawahar et al.
-2023b), and gradient conflicts (Gong and Wang 2022).
-More recently (Jawahar et al.
-2023a), demonstrated that LLMs can serve as effective performance predictors in NAS tasks, providing a promising alternative to traditional methods.
+R&D-Agent Multi-trace exploration R&D-phase Incremental Recombination
+ML-Master MCTS (UCT) with context Incremental
+MLE-STAR Incremental exploration Single-stage Final model ensembling
+[…]
+Table 1: Comparison of LLM-based AutoML systems.
+Scoring Model.
+Our approach of performance scoring is inspired by similar performance prediction methods developed in Neural Architecture Search Elsken et al.
+(2019).
+In NAS, a common approach is to use weight-sharing supernets, where multiple architectures are jointly trained by sharing parameters with a large model, and performance is estimated by evaluating sampled architectures on a validation set Jawahar et al.
+(2023a).
+This method, however, faces challenges such as weight co-adaptation Bender et al.
+(2018), capacity bottlenecks Jawahar et al.
+(2023b), and gradient conflicts Gong & Wang (2022).
+More recently Jawahar et al.
+(2023a), demonstrated that LLMs can serve as effective performance predictors in NAS tasks, providing a promising alternative to traditional methods.
 In contrast to NAS, the AutoML setting typically involves a much broader and less constrained search space, which motivates the exploration of LLM-based performance prediction beyond neural architectures.
-Benchmarks
-Using Kaggle competitions to evaluate autonomous ML systems has become a modern approach in a number of recent benchmarks due to clear metrics, variety of tasks and the ability to compare with human solutions.
-One of the first benchmarks to systematically evaluate autonomous ML agents on Kaggle competitions was MLAgentBench (Huang et al.
-2023), which focused on a small set of tasks with simple baselines, measuring agents’ ability to improve upon them.
-Later, DSBench (Jing et al.
-2024), expanded the scope but often relied on automated filtering, which excluded many complex or non-standard competitions.
-The most recent effort, MLE-Bench (Chan et al.
-2024) stands out for its scale and diversity, present a more challenging and realistic testbed for multi-agent AutoML systems.
+Benchmarks.
+Using Kaggle competitions to evaluate autonomous ML systems has become a modern approach in a number of recent benchmarks due to clear metrics, variety of tasks, and the ability to compare with human solutions.
+One of the first benchmarks for systematically evaluating autonomous ML agents in Kaggle competitions was MLAgentBench Huang et al.
+(2023), which focused on a small set of tasks with simple baselines, measuring agents’ ability to improve on them.
+Later, DSBench Jing et al.
+(2024), expanded the scope but often relied on automated filtering, which excluded many complex or non-standard competitions.
+The most recent effort, MLE-Bench Chan et al.
+(2024) stands out for its scale and diversity, presenting a more challenging and realistic testbed for multi-agent AutoML systems.
 However, MLE-Bench also faces notable limitations, including its large size (3.3 TB) and the fact that it constructs its test sets by partitioning the original training data and then compares the resulting scores to the private leaderboard positions that cause evaluation bias.
@@
 It is designed to ensure robust, leak-free data handling, efficient exploration of modeling ideas, and rapid iteration, all while maintaining high code quality and logical consistency.
-This section outlines the key stages and mechanisms that underpin the operation of our system.
-It consists of four main components: Pipeline Setup, The Ideation Process, Tree-Guided Exploration, which includes operations such as node addition and merging, and the Scoring Model, designed to accelerate overall pipeline evaluation.
-In our framework, each node represents a segment of the final code, and each tree level corresponds to a distinct pipeline component: Exploratory Data Analysis (EDA), Feature Engineering (FE), or Model Training (MT).
-Connections between nodes indicate that they form parts of the same pipeline.
-The main stage unfolds in three phases: tree initialization, adding, and merging.
-Tree initialization constructs the primary tree structure.
-Subsequently, within the allotted time budget, the system alternates between the adding phase — where new ideas are injected into the most promising branches — and the merging phase, which combines two promising ideas into a single, more powerful solution.
-The EDA phase is represented by a single root node, which aggregates all insights derived from data exploration, such as visualizations, distribution analyses, and summary statistics.
-This EDA node is not static: as the search progresses, it can be dynamically enriched with additional analyses prompted by downstream discoveries.
-In contrast, nodes at the FE and MT levels represent concrete instantiations of their respective phases, with edges between nodes indicating their inclusion in the same candidate solution.
-While a node may have multiple children, it is restricted to a single parent.
+This section outlines the key components and mechanisms that underpin the operation of our system.
+It consists of three main stages: (3.1) Pipeline Setup, which prepares the core components; (3.2) Tree Initialization, which generates an initial set of candidate pipelines; and (3.3) Tree-Guided Exploration, the main stage involving tree operations such as node adding and merging, supported by a Scoring Model to speed up pipeline evaluation.
 3.1 Pipeline Setup
+Figure 1: The KompeteAI AutoML pipeline.
 This phase sets up the core components required for the next stages.
@@
 The Metric Agent constructs unit tests to support submission validation and defines the evaluation metric function.
-The Validator Agent partitions the data according to the task specification and applies appropriate preprocessing methods to ensure a valid and reliable evaluation protocol.
+The Validator Agent partitions the data according to the task specification and applies appropriate pre-processing methods to ensure a valid and reliable evaluation protocol.
 The Baseliner Agent generates an initial solution, establishing a lower-bound reference for expected performance.
 Based on the baseline score, it also assesses the quality of the data split and, if necessary, can trigger The Validator Agent to re-partition the data using an alternative strategy.
-3.2 The Ideation Process
+3.2 Tree Initialization
+The primary objective of the tree initialization stage is to generate an initial set of candidate pipelines, providing a diverse foundation from which further exploration can proceed.
+This stage seeds the search space with promising ideas and establishes the initial structure for subsequent optimization.
+3.2.1 Node Representation
+In our framework, each node encodes a code segment, and each tree level maps to a pipeline component: Exploratory Data Analysis (EDA), Feature Engineering (FE), or Model Training (MT).
+Connections between nodes define candidate pipelines.
+The root EDA node aggregates insights from data exploration — visualizations, distributions, and summary statistics — and can be dynamically enriched as the search progresses.
+In contrast, FE and MT nodes represent concrete instantiations of their phases, with edges marking their inclusion in the same solution.
+Each node may have multiple children but only one parent.
+3.2.2 The Ideation Process
 This section focuses on the generation process for a single node within the pipeline, emphasizing localized decision-making rather than full-path construction.
@@
 Finally, the Debugger Agent addresses potential runtime errors and integration issues, completing the node’s initialization cycle.
-Figure 1: The KompeteAI AutoML pipeline consists of three main stages: Pipeline Setup, Tree Initialization, and the Main Loop.
-Pipeline Setup involves data ingestion, initial analysis, metric evaluation, data validation, and establishing performance benchmarks.
-Tree Initialization uses EDA to generate initial insights, forming a tree structure where Feature Engineering nodes are parents to Model Training nodes.
-The Main Loop iteratively refines the solution tree through adding new nodes, with their performance subsequently evaluated via code acceleration, debugging, and return speed to get a score; backpropagation for scoring FE ideas based on downstream model feedback; and merging strong partial solutions to recombine effective strategies
-Insighter
 The Insighter Agent generates ideas for downstream components, guided by factors such as EDA results and the current solution stage.
 To overcome LLM limitations in diversity and quality, it uses two key mechanisms.
-  • Tree Memory: The agent employs a memory module over the entire ideation tree, analyzing previously generated nodes using diversity-driven strategies based on embedding similarity — such as selecting nodes closest or farthest from the parent, or sampling randomly.
+  • Tree Memory: The agent employs a memory module over the entire ideation tree, analyzing previously generated nodes using diversity-driven strategies based on embedding cosine similarity — such as selecting nodes closest or farthest from the parent, or sampling randomly.
 A top-$n$ subset is selected for inclusion in the context, and the memory is continuously updated as new nodes are added, enabling dynamic promotion of diversity.
@@
 The retrieval approach is adaptive, based on function calling that is triggered only when external knowledge is expected to contribute meaningfully, thereby minimizing computational overhead.
-Checker
 The Checker Agent is designed to validate the logical consistency of outputs produced by other agents. It uses unit tests, including schema validation and script execution, to assess correctness.
-Coder
 The Coder Agent implements the input idea based on several parameters, including a description of the input data, available computational resources, and the idea itself.
 It selects optimal hyperparameters and designs the architecture to ensure the code executes within the given time constraints while remaining as efficient as possible.
-Debugger
 The Debugger agent efficiently resolves issues with dependency installation, code generation, and submission formatting using a nested loop that iteratively debugs code within a set limit.
@@
 3.3 Tree-Guided Exploration
-This section provides an explanation of the core principles underlying operations on the ideation tree.
-Tree Initialization
-The primary objective of the tree initialization stage is to generate an initial set of candidate pipelines, providing a diverse foundation from which further exploration can proceed.
-This stage seeds the search space with promising ideas and establishes the initial structure for subsequent optimization.
-After the initial tree is constructed, we perform a backpropagation step to propagate performance signals from the model training nodes (where direct evaluation scores are available) up through the tree, updating each node’s average score based on new information to improve their representativeness within the search tree.
-Adding
+Tree-Guided Exploration has three components: Adding, Merging, and Scoring Model.
+The Tree Initialization stage builds the primary structure.
+Within the time budget, the system alternates between Adding — injecting new ideas into promising branches — and Merging, which combines two strong ideas into a single solution.
+The scoring model works alongside these phases to accelerate pipeline evaluation.
+After each Adding–Merging cycle, we backpropagate performance signals from MT nodes (where evaluation scores are available) upward, updating each node’s average score to improve their representativeness in the search tree.
+The average score of each node is calculated as the mean of the scores of its children.
+3.3.1 Adding
 The adding stage governs the structured expansion of the ideation tree $T_{t}=(V_{t},E_{t})$ by proposing new nodes at the Feature Engineering and Model Training levels.
 Before expansion, the agent examines the current tree state and may selectively trigger additional exploratory analysis via function calling, choosing specific data inspections that are most informative for subsequent node generation.
-The entire process is conditioned on a global context representation $\mathbf{c}_{t}$, which captures the current pipeline state and external knowledge in a structured and denoised form.
-We define the context vector as:
-$\mathbf{c}_{t}=\phi_{\mathrm{EDA}}(T_{t})+\phi_{\mathrm{reader}}(T_{t})+\phi_{\mathrm{ext}}(\texttt{QueryExternal}())$
-where each mapping $\phi:\mathcal{X}\rightarrow\mathcal{C}$ transforms raw, unstructured, or noisy inputs into structured semantic representations in a common context space $\mathcal{C}$.
-Specifically:
-  • $\phi_{\mathrm{EDA}}$ encodes statistical and structural information extracted from exploratory data analysis;
-  • $\phi_{\mathrm{reader}}$ derives contextual insights from competition metadata and prior solutions;
-  • $\phi_{\mathrm{ext}}$ incorporates knowledge retrieved from external sources (e.g., Kaggle, arXiv).
+The process is conditioned on a global context vector $\mathbf{c}_{t}$, representing the current pipeline state and external knowledge in structured form.
+We define $\mathbf{c}_{t}=\phi_{\mathrm{EDA}}(T_{t})+\phi_{\mathrm{reader}}(T_{t})+\phi_{\mathrm{ext}}(\texttt{QueryExternal}())$, where each mapping $\phi:\mathcal{X}\rightarrow\mathcal{C}$ converts noisy inputs into structured semantics in the shared context space $\mathcal{C}$.
+Here, $\phi_{\mathrm{EDA}}$ encodes statistical/structural signals from exploratory analysis, $\phi_{\mathrm{reader}}$ extracts insights from metadata and past solutions, and $\phi_{\mathrm{ext}}$ adds knowledge from external sources such as arXiv or Kaggle.
 Based on the aggregated context $\mathbf{c}_{t}$, the agent performs a structured expansion of the ideation tree in three successive steps.
-First, a set of candidate FE nodes is sampled from the conditional distribution
-$q_{\mathrm{FE}}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{V}_{\mathrm{FE}}),\quad q_{\mathrm{FE}}(\cdot\mid\mathbf{c}_{t})$
-where $\mathcal{V}_{\mathrm{FE}}$ denotes the space of available FE transformations.
-Then, for each sampled FE node $v^{\mathrm{FE}}\in\mathcal{V}_{\mathrm{FE}}$, the agent generates a corresponding set of MT nodes by sampling from
-$q_{\mathrm{MT}}:\mathcal{V}_{\mathrm{FE}}\rightarrow\mathcal{P}(\mathcal{V}_{\mathrm{MT}})$
-which specifies a distribution over the MT configuration space $\mathcal{V}_{\mathrm{MT}}$.
-Finally, a subset of FE nodes from the current tree (including both existing and newly added ones) is selected based on their scores, reflecting their estimated utility.
-For each selected FE node, additional MT nodes are appended as children.
-The full procedure is detailed in Appendix A1.
-Merging
+First, it samples a set of candidate FE nodes according to the context $\mathbf{c}_{t}$, where $\mathcal{V}_{\mathrm{FE}}$ denotes the space of available FE transformations.
+Then, for each sampled FE node $v^{\mathrm{FE}}$, the agent generates a corresponding set of MT nodes within the MT configuration space $\mathcal{V}_{\mathrm{MT}}$.
+Finally, a subset of FE nodes from the current tree (including both existing and newly added ones) is selected based on their scores, and for each selected node additional MT nodes are appended as children.
+The full procedure is detailed in Appendix 1.
+3.3.2 Merging
 The merging stage enables the agent to consolidate multiple promising solutions at the Feature Engineering and Model Training levels, yielding stronger and more generalizable configurations.
@@
     A set of FE node pairs $(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$ is sampled, excluding those present in the long-term memory buffer $\mathcal{M}_{\mathrm{long}}$.
-Each valid pair is merged into a new node
-    $v_{ij}^{\mathrm{FE}}=\texttt{MergeFE}(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$
-    which recombines structural and statistical traits of its parents.
+Each valid pair is merged into a new node $v_{ij}^{\mathrm{FE}}=\texttt{MergeFE}(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$ which recombines structural and statistical traits of its parents.
   2. 2.
@@
 Then, for each parent FE node $v_{i}^{\mathrm{FE}}$ and $v_{j}^{\mathrm{FE}}$, the agent selects additional MT nodes from their respective subtrees.
-These are sampled stochastically with probabilities proportional to their scores:
-    $u_{k}^{(i)}\sim\texttt{SampleTop}(v_{i}^{\mathrm{FE}})$
-    The final child set of $v_{ij}^{\mathrm{FE}}$ combines the freshly generated MT nodes and the resampled top performers from its parents.
+These are sampled stochastically with probabilities proportional to their scores: $u_{k}^{(i)}\sim\texttt{SampleTop}(v_{i}^{\mathrm{FE}})$ The final child set of $v_{ij}^{\mathrm{FE}}$ combines the freshly generated MT nodes and the resampled top performers from its parents.
   3. 3.
-    Additionally, a subset $\mathcal{V}_{\mathrm{FE}}^{\mathrm{merge}}\subseteq V_{t}^{\mathrm{FE}}$ is selected, and within each selected FE node, MT child pairs are merged using
-    $u_{ij}^{\mathrm{MT}}=\texttt{MergeMT}(u_{i},u_{j})$
-    to form stronger model configurations.
+    Additionally, a subset $\mathcal{V}_{\mathrm{FE}}^{\mathrm{merge}}\subseteq V_{t}^{\mathrm{FE}}$ is selected, and within each selected FE node, MT child pairs are merged using $u_{ij}^{\mathrm{MT}}=\texttt{MergeMT}(u_{i},u_{j})$ to form stronger model configurations.
     To avoid redundant or destructive merges, the agent employs dual memory buffers.
@@
 Specifically, a pair that fails to produce a beneficial merge $\theta_{\text{fail}}$ times is promoted from $\mathcal{M}_{\mathrm{short}}$ to $\mathcal{M}_{\mathrm{long}}$, ensuring efficient resource allocation and adaptive learning.
-Merging procedure detailed in Appendix A2.
-3.4 Scoring Model
+Merging procedure detailed in Appendix 2.
+3.3.3 Scoring Model
 The scoring model is a crucial component applied at the model training nodes of our pipeline generation tree.
@@
 This step is essential, as even minor code differences — such as the number of training epochs — can lead to significant variations in model performance.
-These descriptions are then inserted into a prompt to LLM with dataset description and current model training specification.
+These descriptions are then inserted into a prompt to LLM with the dataset description and current model training specification.
 To construct effective anchor ideas for the scoring model, we first train several model architectures on a single feature engineering idea to observe how different architectures behave on the dataset.
@@
 Second, MLE-Bench constructs test sets by partitioning the original training data and compares results to private leaderboard positions.
-This introduces significant evaluation bias, as it does not reflect actual test set performance.
+As shown in table 3, this introduces significant evaluation bias, as it does not reflect actual test set performance.
 We show that scores on MLE-Bench’s constructed test sets can diverge significantly from real leaderboard rankings, leading to misleading conclusions and undermining its reliability as a proxy for real-world outcomes.
@@
   • Selection of Established Competitions.
-We curate 15 competitions from the ’lite’ MLEBench collection that still accept late submissions on Kaggle and whose individual dataset sizes do not exceed 1 GB.
+We curate 15 competitions from the ’lite’ MLE-Bench collection that still accept late submissions on Kaggle and whose individual dataset sizes do not exceed 1 GB.
 The aggregate volume of this subset is 5.3 GB, providing a stable foundation for baseline comparisons.
   • Incorporation of New Competitions.
-To better reflect the evolving landscape of AutoML tasks, we include 11 additional competitions from 2024 and 2025 years.
+To better reflect the evolving landscape of AutoML tasks, we include 11 additional competitions from the years 2024 and 2025.
 These datasets, totaling 4.9 GB, were selected primarily to ensure fair comparison with human performance.
@@
 5.1 Experiment Setup
-Baselines
-To provide a comprehensive evaluation on MLE-Bench, we compared our system against the top-ranked methods on the leaderboard: AIDE, RD-agent, and Ml-Master.
-Due to the high cost of running all baselines, we report their leaderboard metrics as provided directly by MLE-Bench.
-For evaluation on Kompete-bench we re-ran only AIDE and RD-agent, as both are open-source, while the implementation of Ml-Master was not publicly available at the time of our study and was therefore excluded.
-All methods were tested using the same LLM backend — gemini-2.5-flash (Comanici et al.
-2025), which also powers our system.
-Additionally, to fairly assess the alignment between medal distributions on MLE-Bench and real Kaggle leaderboards, we also evaluated all methods using o1-preview (Jaech et al.
-2024), the model on which most MLE-Bench submissions achieved their highest scores.
-Environment
+Baselines.
+To provide a comprehensive evaluation on MLE-Bench, we compared our system with the top open-source methods on the leaderboard: AIDE, RD-agent, ML-Master, and MLE-STAR.
+Due to the high cost of rerunning all baselines, we report their leaderboard metrics as provided by MLE-Bench.
+For Kompete-Bench, all methods were tested with the same LLM backend — gemini-2.5-flash Comanici et al.
+(2025), which also powers our system, except ML-Master, which was run on deepseek-r1 Guo et al.
+(2025) due to architectural constraints.
+To ensure a fair comparison between MLE-Bench and Kompete-Bench and eliminate potential bias from using different LLM backends, we further evaluated AIDE and RD-Agent with o1-preview Jaech et al.
+(2024) — the model on which most MLE-Bench submissions achieved their highest scores.
+Environment.
 To evaluate our system on MLE-Bench, we set a 6-hour runtime limit using 12 vCPUs, 64 GB of RAM, and a single NVIDIA A100 GPU (40 GB).
@@
 Each configuration was executed three times, and results were averaged to reduce variance.
-To ensure fairness and prevent systems from exploiting test-time information, we additionally RAG components from accessing any information about the competition date on which the agent was evaluated.
-Benchmarks
+To ensure fairness and prevent systems from exploiting test-time information, we restrict RAG components in each run from accessing any information about the competition solutions, including those released after its completion.
+Benchmarks.
 We evaluate KompeteAI on the Lite subset of MLE-Bench, a computationally feasible proxy for the full benchmark.
-Although not identical, MLE-Bench reports high alignment in system rankings between Lite and full versions, making Lite a practical basis for comparison.
+Although not identical, MLE-Bench reports high alignment in system rankings between the Lite and full versions, making Lite a practical basis for comparison.
 For medal-based evaluation against real Kaggle leaderboards, we include only Lite competitions that still accept submissions.
 Additionally, we assess performance on our custom benchmark — Kompete-bench, structured following the same principles described in our methodology section, by separating the data into two categories: MLE-subset Lite and recent and a collection of newly curated competitions.
-Evaluation Metrics
+Evaluation Metrics.
 On MLE-Bench, we adopt the official leaderboard metric — the percentage of submissions that receive a medal — which aligns with standard Kaggle competition criteria.
@@
 5.2 Results
-Evaluation on MLE-Bench
+Evaluation on MLE-Bench.
 We begin by evaluating our system on the widely adopted MLE-Bench benchmark (Lite subset).
@@
 ML-Master (deepseek-r1) 48.5 ± 1.5
+MLE-STAR (gemini-2.5-flash) 43.9 ± 6.2
 KompeteAI (gemini-2.5-flash) 51.5 ± 1.5
 Table 2: Comparison of agents on MLE-Bench ’Lite’ subset.
-For AIDE and RD-agent, we use results reported by MLE-Bench.
+For AIDE, RD-agent, MLE-STAR and ML-Master, we use results reported by MLE-Bench.
 KompeteAI was run 3 times with different seeds; results are reported as mean ± SEM.
 Each run was given 6 hours.
-Kaggle-Based Validation of Robustness
+Kaggle-Based Validation of Robustness.
 To validate robustness beyond proxy benchmarks, we tested all agents on the Lite subset using official Kaggle submissions (Table 3).
 The analysis reveals a substantial gap between MLE-Bench evaluations and actual leaderboard outcomes.
-In particular, both AIDE and RD-agent exhibit inflated medal rates on MLE-Bench compared to their real Kaggle performance.
+In particular, all comparable systems exhibit inflated medal rates on MLE-Bench compared to their real Kaggle performance.
 This discrepancy highlights the limitations of relying solely on MLE-Bench for comparison, especially in competitive human-level settings.
-Name MLE-Bench LB Real LB Contemporary part
+Name MLE-Bench LB, % Real LB, % LBs medal intersection, %
 AIDE (o1-preview) 20 7 0
-RD-agent (o1-preview) 27 13 0
+RD-agent (o1-preview) 27 13 7
 AIDE (gemini-2.5-flash) 13 7 0
 RD-agent (gemini-2.5-flash) 13 7 0
-Table 3: Medal rate gap between MLE-Bench leaderboard and leaderboard of real Kaggle competition. For this experiment we ran each system once and used the rules from Kaggle to determine the medal.
-Evaluation on Kompete-Bench
+ML-Master (deepseek-r1) 27 13 7
+MLE-Star (gemini-2.5-flash) 20 13 7
+Table 3: Medal rate gap between the MLE-Bench leaderboard and the leaderboard of a real Kaggle competition.
+The ”LBs medal intersection” column indicates the percentage of competitions where a medal was won on both the real LB and the MLE-Bench LB.
+For this experiment we ran each system once and used the rules from Kaggle to determine the medal.
+Evaluation on Kompete-Bench.
 To address these limitations, Kompete-Bench evaluates performance using real Kaggle leaderboards, covering recent competitions with high prize pools and strong engagement.
@@
 We therefore propose a finer-grained metric: percent humans beaten.
-As shown in Figure 2, KompeteAI demonstrates state-of-the-art results, surpassing human performance in 11.2% of cases — significantly ahead of both RD-agent and AIDE.
+As shown in Figure 2, KompeteAI demonstrates state-of-the-art results, surpassing human performance in 14% of cases — significantly ahead of all comparable systems.
 Nevertheless, the agent remains far behind top leaderboard teams, reflecting how contemporary competitions demand capabilities beyond local modeling — including large-scale feature engineering, synthetic data generation, and use of external datasets.
 Human participants leverage broader tools and resources, underscoring the competitiveness of these challenges and the remaining headroom for automated agents.
-Impact of the Acceleration Methods
+Impact of the Acceleration Methods.
 Finally, we assess the contribution of our acceleration framework.
@@
 While we restrict acceleration measurements to KompeteAI due to tight pipeline integration, the proposed paradigms are general and may be adapted to other architectures with minimal conceptual changes.
-Figure 2: Comparison of our pipeline with AIDE and RD-agent on Contemporary and MLE-Bench parts of Kompete-bench.
-All systems use gemini-2.5-flash as the underlying LLM.
+Figure 2: Comparison of our pipeline with AIDE, RD-agent and ML-Master on Contemporary and MLE-Bench parts of Kompete-bench.
+All systems use gemini-2.5-flash as the underlying LLM, except for ML-Master, which uses deepseek-r1.
 Each was run 3 times with different seeds, results are averaged, and each run was limited to 6 hours.
@@
 […]
-with all accelerations 12.5 ± 2.1
+with all accelerations 12.5 ± 4.1
 Table 4: Impact of acceleration techniques on the number of completed iterations within a fixed time budget. Mean ± 95% CI estimated via Student’s t-distribution.
@@
 6 Conclusions
-We introduced KompeteAI, a new multi-agent AutoML framework, and demonstrated its strong performance across diverse and challenging tasks.
+We introduced KompeteAI, a new multi-agent AutoML framework, and demonstrated its strong performance on diverse and challenging tasks.
+Our contributions include a state-of-the-art architecture with dynamic integration of external knowledge sources and novel tree-based operators for adding and merging, combining LLM decision-making with algorithmic methods.
+We further propose an Accelerated Evaluation and Debugging Paradigm to address execution bottlenecks, applicable to general LLM-based AutoML architectures.
+Finally, we show that the current benchmark suffers from significant bias and introduce Kompete-Bench to overcome this limitation.
+Acknowledgments.
+The research has been supported by the Ministry of Economic Development of the Russian Federation (agreement No. 139-10-2025-034 dd. 19.06.2025, IGK  [^00] C313925P4D0002).
+Appendix A Future work
 Our work opens several promising directions for future research.
@@
 As autonomous agents begin contributing to hypothesis generation, experimental design, and even paper drafting, refining these tools for real research workflows becomes an exciting and urgent frontier.
-Acknowledgements
-The study research has been supported by the Ministry of Economic Development of the Russian Federation (agreement No. 139-10-2025-034 dd. 19.06.2025, IGK  [^00] C313925P4D0002).
-Appendix A Algorithms
-A.1 Adding
+Appendix B Algorithms
+B.1 Adding
 This subsection formalizes the adding stage, as defined in Algorithm 1.
@@
 16:return $T_{t+1}$
-A.2 Merging
+B.2 Merging
 This subsection defines the merging stage and its associated memory mechanism. Let
@@
 25:return $T_{t+1},\mathcal{M}_{\mathrm{short}},\mathcal{M}_{\mathrm{long}}$
-Appendix B Kompete-bench
-Figure A3: End dates for competitions presented in Kompete-Bench
+Appendix C Kompete-bench
+0-99 Teams 100-249 Teams 250-999 Teams 1000+ Teams
+Bronze Top 40% Top 40% Top 100 Top 10%
+Silver Top 20% Top 20% Top 50 Top 5%
+Gold Top 10% Top 10 Top 10 + 0.2%* Top 10 + 0.2%*
+Table 5: Thresholds for winning a medal in Kaggle competitions vary depending on the number of teams participating in each competition.
+We implement the same thresholds as in MLE-bench.
+*the threshold increases by 1 for every 500 additional teams.
+Source: Kaggle (2024).
 Full descriptions of the competitions in Kompete-bench, including the name, number of participants, metrics, and medal thresholds, are listed in Table 5.
@@
 For ongoing competitions, evaluation is performed simultaneously on the public leaderboard.
+Figure A3: End dates for competitions presented in Kompete-Bench
 Name Number of participants Metric Bronze Silver Gold Part aerial-cactus-identification 1221 ROC-AUC ↑ 1 1 1 MLE-bench (Lite) denoising-dirty-documents 162 RMSE ↓ 0.04517 0.02609 0.01794 MLE-bench (Lite) dog-breed-identification 1281 log loss ↓ 0.04598 0.00539 0.0005 MLE-bench (Lite) dogs-vs-cats-redux-kernels-edition 1315 log loss ↓ 0.06127 0.05038 0.03882 MLE-bench (Lite) jigsaw-toxic-comment-classification-challenge 4539 mean col-wise ROC AUC ↑ 0.98639 0.98668 0.98740 MLE-bench (Lite) leaf-classification 1596 log loss ↓ 0.01526 0.00791 0.00000 MLE-bench (Lite) mlsp-2013-birds 81 ROC-AUC ↑ 0.87372 0.90038 0.93527 MLE-bench (Lite) nomad2018-predict-transparent-conductors 879 RMSLE ↓ 0.06582 0.06229 0.05589 MLE-bench (Lite) plant-pathology-2020-fgvc7 1318 ROC-AUC ↑ 0.97361 0.97465 0.97836 MLE-bench (Lite) random-acts-of-pizza 462 ROC-AUC ↑ 0.6921 0.76482 0.97908 MLE-bench (Lite) spooky-author-identification 1242 log loss ↓ 0.29381 0.26996 0.16506 MLE-bench (Lite) tabular-playground-series-dec-2021 1189 ROC-AUC ↑ 0.95658 0.95658 0.9566 MLE-bench (Lite) tabular-playground-series-may-2022 1152 ROC-AUC ↑ 0.99818 0.99822 0.99823 MLE-bench (Lite) text-normalization-challenge-english-language 261 accuracy ↑ 0.99038 0.99135 0.99724 MLE-bench (Lite) text-normalization-challenge-russian-language 163 accuracy ↑ 0.97592 0.98232 0.99012 MLE-bench (Lite) eedi-mining-misconceptions-in-mathematics 1449 MAP@25 ↑ 0.46090 0.49136 0.56429 Contemporary learning-agency-lab-automated-essay-scoring-2 2708 quadratic weighted kappa ↑ 0.83471 0.83518 0.83583 Contemporary lmsys-chatbot-arena 1688 log loss ↓ 1.00472 0.99410 0.98392 Contemporary pii-detection-removal-from-educational-data 2049 efficiency score ↑ 0.95714 0.95883 0.96615 Contemporary um-game-playing-strength-of-mcts-variants 1610 RMSE ↓ 0.43050 0.42973 0.42591 Contemporary llm-prompt-recovery 2176 Sharpened Cosine Similarity ↑ 0.6375 0.6513 0.6848 Contemporary equity-post-HCT-survival-predictions 3327 C-index ↑ 0.69288 0.69320 0.69500 Contemporary cmi-detect-behavior-with-sensor-data 2156 F1 ↑ 0.84 0.84 0.86 Contemporary make-data-count-finding-data-references 833 F1 ↑ 0.548 0.564 0.620 Contemporary neurips-open-polymer-prediction-2025 1539 wMAE ↓ 0.057 0.041 0.032 Contemporary wsdm-cup-multilingual-chatbot-arena 890 categorization accuracy ↑ 0.696381 0.702772 0.711412 Contemporary
 Table 6: Summary of tasks and metrics included in Kompete-bench: the table lists competition names, number of participants, evaluation metrics, threshold values for medals and the part of the benchmark to which this competition belongs
-Appendix C Setup details
+Appendix D Setup details
 We used the hyperparameter settings summarized in Table 8.
@@
 Below are the hyperparameters and their descriptions for each system.
-KompeteAI
-The list of hyperparameters used when running KompeteAI on benchmarks is given in the Table 6.The time_run_minutes parameter sets the maximum runtime for the entire multi-agent system in minutes, after which the process will terminate.
-The runtime_error_time defines the time limit (in minutes) after which a generated code is will be stopped.
+KompeteAI.
+The list of hyperparameters used when running KompeteAI on benchmarks is given in Table 6.The time_run_minutes parameter sets the maximum runtime for the entire multi-agent system in minutes, after which the process will terminate.
+The runtime_error_time defines the time limit (in minutes) after which a generated code will be stopped.
 The subset_size_in_percent specifies the percentage of the dataset to be used for quick validation.
 The validator_size_threshold sets a threshold for the dataset size; if the data exceeds this value, a subset is used for training and validation.
-The number_of_ideas_eda determines how many exploratory data analysis (EDA) ideas generates per iteration.
+The number_of_ideas_eda determines how many exploratory data analysis (EDA) ideas are generated per iteration.
 Similarly, number_of_ideas_data and number_of_ideas_modelling control the number of ideas related to feature engineering and model training.
@@
 The number_of_iterations_parents sets how many iterations parent nodes participate in generating new ideas, while number_of_selected_node_merging determines how many nodes are chosen for merging at each iteration.
-The number_of_ideas_min and number_of_ideas_max define the minimum and maximum number of ideas, which are used as achor ideas for scoring model.
-The retrieve_n_papers and retrieve_n_competitions parameters control how many papers retrieves from arXiv papers and Kaggle solutions.
+The number_of_ideas_min and number_of_ideas_max define the minimum and maximum number of ideas, which are used as anchor ideas for the scoring model.
+The retrieve_n_papers and retrieve_n_competitions parameters control how many papers are retrieved from arXiv papers and Kaggle solutions.
 The number_rag_ideas sets how many ideas are generated using RAG.
@@
 Table 7: Hyperparameters used for running KompeteAI
-RD-Agent
-The list of hyperparameters used when running RD-agent on benchmarks is given in the Table 7.
+RD-Agent.
+The list of hyperparameters used when running RD-agent on benchmarks is given in Table 7.
 The debug_timeout parameter sets the maximum time, in seconds, that is allowed to debug one generated code.
-The full_timeout parameter defines the overall time limit for system.
+The full_timeout parameter defines the overall time limit for the system.
 The if_action_choosing_based_on_UCB flag determines whether the agents select their actions using the Upper Confidence Bound (UCB) strategy.
@@
 Table 8: Hyperparameters used for running RD-agent
-AIDE
-The list of hyperparameters used when running RD-agent on benchmarks is given in the Table 8.
-The steps parameter defines the maximum number of steps the entire system can perform during run.
-The max_debug_depth, specifies the maximum depth for recursive code debugging.
+AIDE.
+The list of hyperparameters used when running RD-agent on benchmarks is given in Table 8.
+The steps parameter defines the maximum number of steps the entire system can perform during a run.
+The max_debug_depth specifies the maximum depth for recursive code debugging.
 The debug_prob parameter suggests that debugging is enabled for every generated code, meaning that all relevant information will be recorded without any sampling.
-Finally, the time_limit parameter represents the maximum allowed wall-clock time (in seconds) for experiment.
+Finally, the time_limit parameter represents the maximum allowed wall-clock time (in seconds) for the experiment.
 Hyperparam name Value steps 2000 max_debug_depth 20 debug_prob 1 time_limit 21600
 Table 9: Hyperparameters used for running AIDE
-Appendix D Metrics
-Competition RD-agent AIDE KompeteAI aerial-cactus-identification 58 5 79 denoising-dirty-documents 3 None 12 dog-breed-identification 41 62 58 dogs-vs-cats-redux-kernels-edition 86 6 91 jigsaw-toxic-comment-classification-challenge 42 26 88 leaf-classification 38 24 46 mlsp-2013-birds None 0 0 nomad2018-predict-transparent-conductors 78 33 22 plant-pathology-2020-fgvc7 12 62 68 random-acts-of-pizza 58 47 67 spooky-author-identification 71 47 59 tabular-playground-series-dec-2021 17 22 25 tabular-playground-series-may-2022 16 32 36 text-normalization-challenge-english-language 3 9 27 text-normalization-challenge-russian-language 17 14 18
+Appendix E Metrics
+Competition RD-agent AIDE ML-Master MLE-STAR KompeteAI aerial-cactus-identification 58 5 58 88 79 denoising-dirty-documents 3 None 15 0 12 dog-breed-identification 37 62 54 44 58 dogs-vs-cats-redux-kernels-edition 86 6 47 84 91 jigsaw-toxic-comment-classification-challenge 40 26 47 30 88 leaf-classification 33 24 35 54 46 mlsp-2013-birds None 0 9 0 None nomad2018-predict-transparent-conductors 78 33 76 30 22 plant-pathology-2020-fgvc7 12 62 45 38 68 random-acts-of-pizza 58 47 56 58 70 spooky-author-identification 67 47 60 58 63 tabular-playground-series-dec-2021 17 22 26 13 29 tabular-playground-series-may-2022 16 32 29 23 43 text-normalization-challenge-english-language 3 9 25 7 27 text-normalization-challenge-russian-language 17 14 36 33 18
 Table 10: Full table by ”percent humans beaten” for AutoML systems given in the article at each competition in MLE-Bench part of Kompete-bench
-Competition RD-agent AIDE KompeteAI cmi-detect-behavior-with-sensor-data None None None eedi-mining-misconceptions-in-mathematics 22 None 30 equity-post-HCT-survival-predictions None None None learning-agency-lab-automated-essay-scoring-2 12 19 13 llm-prompt-recovery 12 9 None lmsys-chatbot-arena None None None make-data-count-finding-data-references 11 None 18 neurips-open-polymer-prediction-2025 13 None 15 pii-detection-removal-from-educational-data None None 29 um-game-playing-strength-of-mcts-variants None None None wsdm-cup-multilingual-chatbot-arena 14 None 19
+Competition RD-agent AIDE ML-Master MLE-STAR KompeteAI cmi-detect-behavior-with-sensor-data None None None None None eedi-mining-misconceptions-in-mathematics 19 None 18 9 30 equity-post-HCT-survival-predictions None None None None None learning-agency-lab-automated-essay-scoring-2 9 23 15 17 13 llm-prompt-recovery 11 11 17 None None lmsys-chatbot-arena None None 14 17 15 make-data-count-finding-data-references 13 None 17 21 18 neurips-open-polymer-prediction-2025 9 None 10 11 21 pii-detection-removal-from-educational-data None None 19 13 38 um-game-playing-strength-of-mcts-variants None None None None None wsdm-cup-multilingual-chatbot-arena 14 None 8 10 19
 Table 11: Full table by ”percent humans beaten” for AutoML systems given in the article at each competition in Contemporary part of Kompete-bench.
-Results of AutoML system on Kompete-bench are presented in the Table 9 and Table 10.
+Results of the AutoML system on Kompete-bench are presented in Table 9 and Table 10.
 We report results using the ”percent humans beaten” metric, which reflects the percentage of Kaggle leaderboard participants outperformed by each system.
@@
 Notably, the primary factor influencing overall performance was the proportion of valid submissions: systems that consistently generated correct code achieved substantially higher scores.
+Appendix F Analysis
+F.1 Performance Assessment of the Scoring Model
+Figure A4: Comparison between predicted scores and actual validation scores after min–max normalization within each competition.
+Each point corresponds to a solution from a distinct competition.
+The solid black diagonal represents the ideal case where predicted and actual scores are perfectly aligned.
+To evaluate the proposed scoring model, we conducted experiments across multiple benchmark competitions.
+Figure A4 compares predicted performance estimates with actual scaled scores obtained from validation splits of the data, covering tasks from image recognition to NLP and structured data.
+To account for the fact that competitions differ substantially in their metric ranges, we normalize scores within each competition using min–max scaling:
+$s^{\text{scaled}}_{i}=\frac{s^{\text{val}}_{i}-\min_{j}s^{\text{val}}_{j}}{\max_{j}s^{\text{val}}_{j}-\min_{j}s^{\text{val}}_{j}},$
+where $s^{\text{val}}_{i}$ denotes the validation score of solution $i$ within a competition.
+This adjustment ensures that relative differences are preserved while eliminating distortions caused by heterogeneous metric ranges, which would otherwise make cross-task comparisons less interpretable.
+The results show a clear positive correlation between predicted and real scores.
+The model not only distinguishes strong solutions from weaker ones but also preserves relative ordering across a wide range of performance levels, which is critical for ranking-based exploration.
+At the same time, deviations from the diagonal indicate systematic tendencies: predictions for top-performing submissions are slightly conservative, while weaker candidates are occasionally overestimated.
+This asymmetry suggests that the model captures the general performance landscape but remains imperfect in calibrating extreme cases.
+Importantly, such biases are less problematic for the intended use case, where relative ranking — rather than precise absolute accuracy — guides the search process.
+Two caveats remain: prediction errors may accumulate during long-running searches, and the evaluation is restricted to Kaggle-style benchmarks.
+Extending validation to real scientific discovery tasks represents an important future direction.
+F.2 Impact of Competition and Literature Retrieval in RAG
+Configuration Variant Relative Performance (%)
+Without competition knowledge 86.7
+Without arXiv knowledge 92.2
+Without any RAG input 83.4
+Table 12: Ablation of retrieval-augmented knowledge sources on a subset of benchmark competitions.
+Reported values indicate the relative performance by competition metric of each ablated configuration compared to the complete RAG-enabled system (treated as the 100% reference point).
+The ablation analysis in Table 12 shows how different retrieval sources contribute to overall system performance.
+Removing competition data reduces relative accuracy to 86.7%, suggesting that competition write-ups often contain concrete heuristics and task-specific tricks that transfer effectively to our benchmark problems.
+Excluding arXiv papers has a milder effect (92.2%).
+This is consistent with the idea that publications tend to capture broader methodological advances.
+While such knowledge is not always directly applicable to competition-style tasks, it still provides useful algorithmic patterns that enhance performance.
+The largest drop is observed when retrieval augmentation is disabled entirely (83.4%).
+This indicates that external knowledge, regardless of source, is a key component of the approach.
+Together, competition reports and academic papers complement each other: the former supply practical, domain-oriented techniques, while the latter contribute generalizable methods and perspectives.
+F.3 Distribution of Successfully Integrated Research Ideas
+Figure A5: Average outcomes of the debugging process.
+The chart illustrates the proportion of successfully integrated ideas, categorized by their target AutoML stage — Model Training (MT) or Feature Engineering (FE) — versus the proportion of ideas that were ultimately discarded as ”Failed”.
+Figure A5 summarizes the outcomes of the debugging and integration pipeline.
+Overall, 80.4% of proposed ideas were successfully incorporated into the research tree, confirming the reliability of debugging mechanisms.
+Only 19.6% of ideas failed to pass validation, that there are significantly fewer ideas that our system successfully debug.